A. Bugzy Company had cash sales of 300,000.00 and credit sales of 200,000.00. Its cost of sales is One Hundred Twenty Thousand Pesos. Accounts Receivable beginning and end balances are 45,000.00 and 20,000.00 respectively. Merchandise Inventory is valued at 100,000.00 on January 1 and 115,000.00 on Devember 31. Answer the questions that follow:



1. What is the company's gross profit margin? How did you compute for such?



2. How many days can you collect the receivables of the company in 2020? How were you able to solve for it?

1
Expert's answer
2021-11-15T12:16:10-0500

Solution:

1.). Gross profit margin:

Gross profit margin = Total sales – Cost of sales = 500,000 – 120,000 = 380,000

Gross profit margin = 380,000

 

2.). The formula for days sales outstanding = (Average accounts receivable "\div" Credit sales) "\times" 365

 

Average accounts receivable = (45,000 + 20,000)"\div" 2 = 32,500

Credit sales = 200,000

DSO = 32,500"\div" 200,000 "\times" 365 = 59 days


To derive days to collect receivables, divide the average accounts receivable during a given period by the total value of credit sales during the same period and multiply the result by the number of days in the period being measured.